# On the Geometry of Semantics in Next-token Prediction

### v1

@@ Abstract
Modern language models demonstrate a remarkable ability to capture linguistic meaning despite being trained solely through next-token prediction (NTP). We investigate how this conceptually simple training objective leads models to extract and encode latent semantic and grammatical concepts. Our analysis reveals that NTP optimization implicitly guides models to encode concepts via singular value decomposition (SVD) factors of a centered data-sparsity matrix that captures next-word co-occurrence patterns. While the model never explicitly constructs this matrix, learned word and context embeddings effectively factor it to capture linguistic structure. We find that the most important SVD factors are learned first during training, motivating using spectral clustering of embeddings to identify human-interpretable semantics, including both classical k-means and a new orthant-based method directly motivated by our interpretation of concepts. Overall, our work bridges distributional semantics, neural collapse geometry, and neural network training dynamics, providing insights into how NTP’s implicit biases shape the emergence of meaning representations in language models.
@@ 1 Introduction
Next-token prediction (NTP) is conceptually simple: predict the next word given a preceding sequence (aka context). Yet it trains models with remarkable ability to capture meaning.
A foundational step toward understanding how NTP leads to models acquiring semantic information is to first understand how explicit textual inputs—words and contexts—are represented in the models’ $d$-dimensional word and context vector representations (embeddings). Zhao et al. [2024] recently proved that in well-trained models, the geometry of these learned representations is characterized by the factorization of a data-sparsity matrix $\widetilde{\bm{S}}$ whose entries encode whether a particular word follows a given context in the training corpus.
While this characterization explains how models encode explicit training signals, it does not address the deeper question about semantic learning. Language conveys meaning through latent concepts that exist beyond explicit (context, word) pairs, leading us to investigate: How do NTP-trained models extract and encode these latent concepts? For instance, how do they capture semantic dichotomies like male/female, or grammatical categories like nouns and verbs? In other words, how does latent information shape the embedding geometry, and does this suggest ways to identify interpretable semantic structures in the embedding space? These questions are challenging because linguistic concepts are never explicitly given as inputs in the NTP objective. Yet, by analyzing the geometry of learned representations and its relation to the data-sparsity matrix, we show that latent linguistic structure emerges from the latter. Concretely, our contributions are summarized as follows.
Geometric emergence of latent concepts: We demonstrate that latent linguistic concepts learned by a large neural model through NTP emerge as principal components in the singular value decomposition of a centered data-sparsity matrix $\tilde{\mathbf{S}}$ (defn. in Sec. 3.1, see Fig. 1 for illustration). While $\widetilde{\bm{S}}$ is never explicitly formed during training, its structure naturally shapes the geometry of learned representations: concepts can be recovered as weighted combinations of word and context embeddings, where the weights are determined by the singular vectors of $\tilde{\mathbf{S}}$. This extends the characterization of the unconstrained-features model (UFM) by Zhao et al. [2024], revealing how embeddings inherently organize around latent concept dimensions.
Rate of learning: Singular values of the data-sparsity matrix quantify the significance of their corresponding concepts. By recognizing a connection between UFM with square-loss and the closed-form training dynamics of two-layer neural networks with orthogonal inputs derived by Saxe et al. [2013], we show that concepts associated with larger singular values are learned faster during training.
Semantics [^1] We broadly define “semantic” to encompass linguistic factors that humans consider when producing text, including grammatical (e.g., tense), syntactic (e.g., verb), and semantic (e.g., sentiment) elements. as orthant-based clustering: Not all concepts have human-interpretable meaning. Guided by the insight that top concepts are learned first during training, we find that human-interpretable semantic information is captured through specific positive/negative configurations (indicating how present/absent the specific concept is in that word or context) across a few top singular concept dimensions (see Fig. 2B for illustration). Identifying semantics through such sign configurations geometrically manifests as (spectral) clustering of word/context embeddings in orthant slices. Through experiments, we verify that this approach identifies distinct human-interpretable linguistic categories.
(A) Left Singular Matrix ($\bm{U}$)Refer to caption(B) Singular ValuesRefer to caption(C) Right Singular Matrix (${\bm{V}}$)Refer to caption
Figure 1: Illustration of semantic information in SVD of $\widetilde{\bm{S}}$ (rank=6). (A) Left singular matrix (U) where rows are analyzer vectors and columns are singular dimensions. The first dimension distinguishes between plants (“apple”, “tulip”, negative values) and animals (“dog”, “shark”, positive values). The second dimension further differentiates mammals (“dog”, “cat”, positive) from other animals (“trout”, “eagle”, negative). (B) Singular values in descending order of importance. (C) Right singular matrix (V) where the first dimension separates plant properties (“uses sunlight for food”) from animal properties (“hunts in groups”). The second dimension appears more complex: for animal properties, it contrasts mammal-specific traits (“stays warm-blooded”, positive) against aquatic traits (“lives in water”, negative), while for plant properties, it differentiates growth characteristics (“grows roots”, positive) from reproductive features (“reproduces with pollen”, negative).
@@
Figure 2: NTP-trained word embeddings naturally organize into semantic clusters. (C) Training data (showing ${\bm{P}}^{\top}$) consists of pairs using the template “The organism that [attribute] is [subject]”, where attributes are rows and subjects are columns. (A) When trained with $d>V>rank(\widetilde{\bm{S}})=6$ and visualized in 3D PCA projection, both words (dots) and contexts (triangular marks) cluster by semantic category—animals (green/blue), and plants (red/orange). This demonstrates how semantic structure emerges naturally from the underlying data sparsity pattern, a process our work formalizes mathematically. (B) Visualization of orthant-based clustering and hierarchical semantic structure in a 3D subspace of context-analyzer vectors. Selected dimensions $\mathcal{K}=\{k_{1},k_{2},k_{3}\}={1,2,4}$ correspond to columns of ${\bm{V}}$ in Fig. 1C. Colored blocks represent different sign configurations: Green ($p=1$): $c_{k_{1}}=-1$ captures "animal" semantic. Blue ($p=2$): $[c_{k_{1}},c_{k_{2}}]=[-1,-1]$ refines to "mammal" semantic. Yellow ($p=3$): $[c_{k_{1}},c_{k_{2}},c_{k_{3}}]=[-1,-1,-1]$ further specifies "feline" semantic. This nested structure illustrates how increasing the number of active dimensions ($p$) yields progressively finer-grained semantic categories, where "mammal" is a subset of "animal", and "feline" is more specific within "mammal". Contexts shown with same color within each suborthant are example members according to Defn. 1.
Connecting distributional semantics and neural collapse (NC) geometries: On the one hand, linking to classical distributional semantics approaches based on heuristic forms of matrix factorization of data co-occurrences, we find using model abstractions popularized in the NC literature that NTP implicitly favors SVD factorization of the centered data-sparsity matrix. On the other hand, we broaden the scope of NC geometries of embeddings to encompass geometries of latent semantics – making the connection explicit, we show the emergence of semantics even for the default one-hot data setting in the NC literature.
2 Related Work
We bring together, in the context of NTP-trained language models, the following two classical ideas: (1) semantic patterns emerge from linear relationships between vector representations of textual inputs, and (2) these learned representations form geometries tied to language statistics via count matrices.
Semantics from Distributed Representations. The idea of extracting semantics from distributed representations, while rooted in earlier works Bengio et al. [2000], was popularized through word embedding research Mikolov et al. [2013a]. Two classically popular approaches include: (i) constructing semantic axes from centroids of pole words An et al. [2018], Fast et al. [2016], evolving from lexicon-based sentiment analysis Taboada et al. [2011], Hu and Liu [2004], and (ii) investigating semantic patterns through “word analogies,” which reveal that semantic relationships manifest as linear directions in embedding space Mikolov et al. [2013c], Levy et al. [2015], Drozd et al. [2016], Ethayarajh et al. [2019], Allen and Hospedales [2019]. In the context of modern LLMs, word analogies have been empirically observed Rezaee and Camacho-Collados [2022], Wijesiriwardene et al. [2024]. However, recent focus has shifted to techniques for distilling interpretable semantics through dictionary learning or sparse autoencoders Cunningham et al. [2023], Bricken et al. [2023], further showing these can be used for model steering Turner et al. [2023], Li et al. [2024] (techniques that in fact trace back to early work with static word embeddings, e.g., Murphy et al. [2012], Faruqui et al. [2015], Arora et al. [2018], Zhang et al. [2019]). Alternative approaches include geometric studies showing how linguistic features cluster in embedding spaces Coenen et al. [2019], Hewitt and Liang [2019]. Simultaneously, supervised methods such as linear probing have been employed to extract semantic and syntactic information from modern LLMs Alain and Bengio [2017], Hewitt and Manning [2019], Marks and Tegmark [2023], relying on the same principle of linear representation found in word analogies. Unlike these aforementioned works that empirically analyze trained models post-hoc, we focus on formalizing the mechanisms by which semantic structure emerges during NTP training. Our goal is not to introduce a state-of-the-art method for extracting semantics, but rather to understand, from an optimization viewpoint, how semantics emerge as a consequence of next-token prediction. While Park et al. [2023, 2024] have explored similar questions, here, rather than relying on a model for concept generation and its link to words/contexts, we directly analyze NTP.
Geometry from Co-occurrence Statistics. Vector space representations emerge from two approaches: (i) traditional count-based models using co-occurrence matrices, and (ii) prediction-based models that optimize token prediction objectives like Skip-gram or NTP. Levy and Goldberg [2014], Pennington et al. [2014] demonstrate that these approaches are fundamentally similar, as both probe underlying corpus co-occurrence statistics. Zhao et al. [2024] have extended this insight to causal NTP models with a key distinction: the co-occurrence matrix between contexts and words is inherently sparse, unlike the dense word-word co-occurrences in prior work. This distinction makes relevant recent advances in implicit regularization Ji and Telgarsky [2019], Ji et al. [2020] and neural collapse geometries Papyan et al. [2020b], Mixon et al. [2020], Fang et al. [2021], through which Zhao et al. [2024] analyze how word and context embeddings converge via an unconstrained features model, relating them to singular factors of the centered co-occurrence probability matrix’s support set (see Eq. (2)). We show that these singular factors encode rich semantic structure. While we argue that the natural language’s inherent multilabel nature is responsible for the emergence of rich semantics, making full circle to the neural-collapse literature, we show that the semantic interpretation applies to as simple a setting as one-hot imbalanced classification. Moreover, by connecting to the closed-form analysis of learning dynamics of linear neural networks by Saxe et al. [2013, 2019], Gidel et al. [2019], we show how singular values determine the learning order during training.
More details on related works deferred to App. E. See also Remark 1.
3 Background
Notations. For any integer $k$, $[k]:=\{1,\ldots,k\}$. Matrices, vectors, and scalars are denoted by $\bm{A}$, $\bm{a}$, and $a$ respectively. For matrix $\bm{A}$, $\bm{A}[i,j]$ is its $(i,j)$-th entry, and for vector $\bm{a}$, $\bm{a}[i]$ is its $i$-th entry. $\mathds{1}$ is the all-ones vector and $\otimes$ is Kronecker product.
3.1 NTP Objective as Sparse Soft-Label Classification

### v2

# Geometry of Semantics in Next-Token Prediction: How Optimization Implicitly Organizes Linguistic Representations

@@ Abstract
We investigate how next-token prediction (NTP) optimization leads language models to extract and organize semantic structure from text. Our analysis, based on a tractable mathematical model and controlled synthetic data, reveals that NTP implicitly guides models to factor a centered support matrix encoding context-to-next-token co-occurrence patterns via singular value decomposition (SVD). While models never explicitly construct this matrix, learned word and context embeddings converge to its SVD factors, with singular vectors encoding latent semantic concepts through their sign patterns. We demonstrate that concepts corresponding to larger singular values are learned earlier during training, yielding a natural semantic hierarchy where broad categories emerge before fine-grained ones. This insight motivates orthant-based clustering, a method that combines concept signs to identify interpretable semantic categories. We validate our findings on synthetic datasets and pretrained language models, recovering diverse semantic structures such as grammatical categories, named entity types, and topical distinctions (medical, entertainment). Our work bridges classical distributional semantics and neural collapse geometry, characterizing how gradient-based optimization implicitly determines both the matrix representation and factorization method that encode semantic structure.
@@ 1 Introduction
Models trained with next-token prediction (NTP) show a remarkable ability to capture and represent linguistic meaning. How does this semantic ability emerge during training? A primary challenge in addressing this question lies in quantifying “meaning”. We therefore focus specifically on understanding the semantic organization of learned representations and the mechanisms by which it arises. Our goal is fundamentally different from the substantial recent interest in post-hoc evaluation of interpretable semantics in Large Language Models (LLMs). Techniques like dictionary learning and sparse autoencoders for model steering Cunningham et al. [2023], Bricken et al. [2023], Turner et al. [2023], Li et al. [2024]—tracing back to similar work with static word embeddings for word2vec architectures Murphy et al. [2012], Faruqui et al. [2015], Arora et al. [2018], Zhang et al. [2019]—aim to extract interpretable features from trained models. In contrast, we seek to understand the specific mechanisms by which gradient-descent-based training, in conjunction with the structure of textual input data, guides a language model to organize its representations into semantically interpretable structures.
Our starting point is the recent work by Zhao et al. [2024], who characterized the geometry of the model’s explicit textual input representations, i.e., $d$-dimensional word/context embeddings. At convergence, this geometry is precisely characterized by the factorization of a specific data-sparsity matrix $\widetilde{\bm{S}}$, which encodes co-occurrence patterns in the training corpus: its entries indicate whether a particular word follows a given context.
This characterization, however, leaves open the question: How do these textual input patterns impact the geometry of implicit linguistic information, such as grammatical or semantic categories, through the optimization of the NTP objective? Specifically, we ask three subquestions:
Since semantic information is not explicitly labeled in the data,
Q1. what latent signals implicitly guide NTP optimization to specific structures?
Q2. How are these structures related to human-interpretable semantics?
Q3. How fast do these semantic structures emerge during training?
Figure 1: Illustration of C1-3. Details in Sec. 3. (A) SVD of the data-sparsity matrix $\widetilde{\bm{S}}$ (Eq. (2)), derived from the training data in Fig. 2(C). Columns of $\bm{U}$ and $\bm{V}$}, serve as word and context concept basis vectors, respectively: Their entries encode an association score with sign indicating the direction (association/opposition) between word/contexts and concepts. The singular values indicate the strength of each concept. (B) Combining a few top concepts according to the specific sign patterns of concept basis, which we call orthant-based clustering, reveals human-interpretable semantics. (C) Gradient-descent-based NTP optimization leads to concepts corresponding to larger singular values being learned first. This corresponds to the emergence of coarse semantics before fine-grained ones, which we empirically validate with visualizations of the trained model’s output confusion matrices.
Towards addressing these questions, we adopt a bottom-up, theory-guided approach. We build our insights by analyzing the optimization of an abstract mathematical model, which isolates the essential components of NTP optimization in a tractable framework (Sec. 2). To validate and further shape these insights, we carefully construct synthetic data settings and formally verify our findings, which can otherwise be difficult to precisely quantify in complex, real-world datasets (Sec. 3). Finally, we demonstrate that the insights gained extend qualitatively to real data and models (Sec. 4). Concretely, our contributions are:
C1. Interpreting latent signals (Sec. 3.1): We show that the singular vectors of $\tilde{\mathbf{S}}$ encode latent concepts: the sign of each entry indicates whether a word or context supports or opposes a concept, and the magnitude reflects the strength of association. This equips the abstract geometry in Zhao et al. [2024] with semantic structure. Fig. 1A visualizes semantics aligning with concept signs.
C2. Establishing human-interpretable semantic structure (Sec. 3.2): Building on C1, we introduce an orthant-based clustering method that identifies interpretable semantic categories by combining a few top concepts according to their sign patterns. We empirically demonstrate that this method recovers a wide array of interpretable linguistic structures—such as part of speech, verb tense, and named entity types—on both synthetic data and real datasets like TinyStories and WikiText-2, where the data-sparsity matrix is explicitly computable. Furthermore, we show in GPT-2, BERT, and Qwen that increasing the number of singular components systematically reveals finer-grained semantics. Fig. 1B visualizes a hierarchy of interpretable semantics via orthant-based clustering.
C3. Analyzing semantic emergence (Sec. 3.3): Beyond convergence, we analyze semantic emergence during training. Theoretically, we link NTP-UFM gradient descent dynamics to the framework of Saxe et al. [2013], showing that top concepts, those corresponding to larger singular values, are learned earlier. Fig. 1C visualizes this phenomenon: broad semantics corresponding to top concepts are formed early during training.
These contributions bridge classical distributional semantics (e.g., Ding et al. [2006], Levy et al. [2015]) and recent neural collapse (NC) geometries Papyan et al. [2020a], Zhao et al. [2024]. By employing tractable abstractions from the latter, we provide theoretical grounding for the former: we show that gradient-based optimization implicitly determines both the matrix representation (data-sparsity matrix) and factorization method (SVD) for encoding semantics, rather than requiring hand-engineered choices (Sec. 5). Conversely, we extend the scope of NC geometries beyond simple classification settings to characterize the underlying hierarchical semantic representations.
Notations. For any integer $k$, $k]:=\{1,\ldots,k\}$. Matrices, vectors, and scalars are denoted by $\bm{A}$, $\bm{a}$, and $a$, respectively. For matrix $\bm{A}$, $\bm{A}[i,j]$ is its $(i,j)$-th entry, and for vector $\bm{a}$, $\bm{a}[i]$ is its $i$-th entry. $\mathds{1}$ is the all-ones vector and $\mathds{I}$ is the identity matrix.
2 Background: Geometry of Word and Context Representations
The NTP objective minimizes the cross-entropy (CE) loss between the model’s output distribution given a context—a sequence $\bm{z}$ of tokens from vocabulary of size VV—and the one-hot representation of the next-word/token. Since multiple tokens can follow a given distinct context $\bm{z}_{j}$ in a dataset, this is equivalent to minimizing CE between the model’s output distribution and an associated soft-label vector $\hat{\bm{p}}_{j}$ representing the conditional frequency distribution of next tokens for that specific context. Because not all vocabulary tokens are valid next-tokens for a given context (even when ignoring the finite size of the data), the soft-label vectors $\hat{\bm{p}}_{j}$ are sparse. For a dataset with $m$ distinct contexts, define the support matrix $\bm{S}\in\{0,1\}^{V\times m}$ of the soft-label matrix ${\bm{P}}=[\hat{\bm{p}}_{1},\ldots,\hat{\bm{p}}_{m}]\in\mathbb{R}^{V\times m}$, such that $\bm{S}[z,j]=1\,\Leftrightarrow\,{\bm{P}}[z,j]>0$.
Building on this formulation, Zhao et al. [2024] investigated how the soft-label structure shapes the learned geometry when optimizing the NTP objective with gradient descent. Specifically, they studied word embeddings (the rows of the $V\times d$ decoding matrix $\bm{W}$}) and context embeddings (the deep $d$-dimensional representations $\{h(\bm{z}_{j})\}_{j\in[m]}$ of distinct contexts) of a model at convergence.
For tractability, they relied on two modeling simplifications: (1) using the limit of vanishing ridge regularization as a proxy for the convergence limit of gradient-descent iterations, and (2) assuming a sufficiently expressive neural network that optimizes context embeddings freely, instead of abiding by an architecture-specific parameterization. The first simplification is standard in the implicit-bias optimization literature Ji et al. [2020], Wei et al. [2019], Tarzanagh et al. [2023], while the second has been used for language models by Yang et al. [2017] and recently popularized in the neural collapse literature Mixon et al. [2022], Wojtowytsch [2021], Fang et al. [2021].
Taken together, these abstractions yield a tractable model for NTP optimization, which we refer to as the unconstrained features model (UFM) for NTP training, or NTP-UFM in short: [^1] Note the structural similarity of this log-bilinear objective to word2vec architectures Mikolov et al. [2013c], Pennington et al. [2014]. However, NTP-UFM serves as a tractable analytical abstraction rather than a direct practical architecture.
$\min_{{\bm{W}},{\bm{H}}}\mathcal{L}({\bm{W}}{\bm{H}};{\bm{P}})+{{\lambda}}\|{\bm{W}}\|^{2}+{{\lambda}}\|{\bm{H}}\|^{2}\,.$ (1)
NTP-UFM jointly optimizes the matrices of word embeddings ${\bm{W}}\in\mathbb{R}^{V\times d}$ and context embeddings ${\bm{H}}:=[\bm{h}_{1},\ldots,\bm{h}_{m}]\in\mathbb{R}^{d\times m}$. Here, $\bm{h}_{j}$ represents $h(\bm{z}_{j})$ and is freely optimized for each context; $\lambda$} is the regularization parameter; and $\mathcal{L}$ is the CE loss acting on the logit matrix $\bm{L}={\bm{W}}{\bm{H}}$, parameterized by the soft-label matrix $\bm{P}$}. Analyzing NTP-UFM, for large embedding dimensions $d\geq V$,  [^2] This does not restrict the number of contexts ($m$, potentially $\gg d$), enabling the study of rich semantic catgories (Sec. 3). Also, qualitative insights extend to real models (Sec. 4.2). and $\lambda\rightarrow 0$, Zhao et al. [2024] show that:
1. Logits: The logit matrix $\bm{L}$ grows unboundedly during training. After normalization, it converges in direction to a matrix $\bm{L}^{\rm{mm}}$ that can be explicitly computed given the data support matrix $\bm{S}$.
2. Word/Context Embeddings: Like logits, word and context embeddings grow unboundedly in magnitude, but converge in direction to $\bm{U}\sqrt{\bm{\Sigma}}\mathbf{R}$ and $\mathbf{R}^{\top}\sqrt{\bm{\Sigma}}{\bm{V}}^{\top}$, respectively. Here, $\bm{U}\bm{\Sigma}{\bm{V}}^{\top}$ is the SVD of $\bm{L}^{\rm{mm}}$ and $\mathbf{R}$ is a partial orthogonal matrix.
3. Data-sparsity matrix as proxy: An accurate proxy for $\bm{L}^{\rm{mm}}$ that is very easy to compute is the centered support matrix
$\widetilde{\bm{S}}:=(\mathds{I}_{V}-\nicefrac{{1}}{{V}}\cdot\mathds{1}_{V}\mathds{1}_{V}^{\top})\,\bm{S}\,$ (2)
$\widetilde{\bm{S}}[z,j]=\begin{cases}1-\frac{1}{V}\sum_{z\in[V]}\bm{S}[z,j]&\text{if $\bm{S}[z,j]=1$}\\ -\frac{1}{V}\sum_{z\in[V]}\bm{S}[z,j]&\text{if $\bm{S}[z,j]=0$}\end{cases}\,$,
which we refer to as the data-sparsity matrix. Note that unlike the support matrix $\bm{S}$, the entries of $\widetilde{\bm{S}}$ are not binary due to the centering operation.
Taken together, Zhao et al. [2024] establishes how textual input structure maps through NTP optimization to word and context representation geometry. Textual input structure is summarized through the SVD of the data-sparsity matrix, where $\bm{U}\in\mathbb{R}^{V\times r}$, ${\bm{V}}\in\mathbb{R}^{m\times r}$ with $\bm{U}^{\top}\bm{U}={\bm{V}}^{\top}{\bm{V}}=\mathds{I}_{r}$, singular values $\bm{\Sigma}=\operatorname{diag}(\sigma_{1},\ldots,\sigma_{r})$ are ordered as $\sigma_{1}\geq\ldots\geq\sigma_{r}>0$, and rank $r\leq V-1$:
$\widetilde{\bm{S}}=\bm{U}\bm{\Sigma}{\bm{V}}^{\top}.$ (3)
The geometry of learned representations is characterized by word and context embeddings that converge in the direction to $\bm{U}\sqrt{\bm{\Sigma}}\mathbf{R}$ and $\mathbf{R}^{\top}\sqrt{\bm{\Sigma}}{\bm{V}}^{\top}$, respectively, with $\mathbf{R}$ a partial orthogonal matrix.
@@ 2 Background: Geometry of Word and Context Representations
Figure 2: Optimizing NTP on (context,next-word) input pairs yields a specific geometry of representation of those inputs; orthant-based clustering reveals how this geometry is organized on the basis of interpretable semantic categories. (A) Empirical 3D PCA visualization of learned word(dots)/context(triangular marks) embeddings from two-layer transformer trained on the dataset in (C) reveals a clear clustering by semantic category (plants in green/blue, animals in red/orange), despite no explicit semantic signal at training. (B) We show (Sec. 3.2) that semantic categories are organized into a hierarchy of sub-orthants defined by concepts combined with respect to their sign patters. For example, Felines occupy a 2D sub-orthant (combining the second and fourth concepts), while Mammals and Animals occupy progressively larger sub-orthants. (C) The sparse training data matrix (rows = attributes, columns = subjects) follows the template: “The organism that [attribute] is [subject].”
3 Geometry of Semantics
The previous section establishes that word and context embeddings converge to the SVD factors of the data-sparsity matrix $\widetilde{\bm{S}}$. But: (i) What latent signals implicitly guide optimization to produce semantic structures within this geometry? (ii) How are these related to human-interpretable semantics? (iii) How fast do they emerge during training?
To motivate our analysis, consider Fig. 2A visualizing word/context embeddings of a transformer trained on (sequence,next-token) pairs from the sparse soft-label matrix $\bm{P}$} shown in Fig. 2C. Despite no explicit semantic supervision during training, clear distinctions emerge. For example, animal-related words and contexts cluster on the left, while plant-related ones cluster on the right. What drives such an implicit semantic structure?
3.1 Identifying Concepts as Latent Signals
Our key insight is to examine the structure of the SVD factors of the data-sparsity matrix $\widetilde{\bm{S}}$. Recall word embeddings take the form ${\bm{W}}=\bm{U}\sqrt{\bm{\Sigma}}\mathbf{R}$, where $\bm{U}$ encodes the row-space of $\widetilde{\bm{S}}$. Since rows of $\widetilde{\bm{S}}$ (and, thus, of $\bm{U}$) correspond to different words, their inner product captures word similarity: words appearing in similar contexts have high similarity. However, semantic relationships extend beyond simple pairwise similarity or synonymy.
To reveal such semantic structure, we instead focus on the columns of $\bm{U}$, which correspond to $r$ latent features in the representation space. We term each SVD dimension $k\in[r]$ a concept. Adopting terminology from Saxe et al. [2019], we call the columns $\mathbf{u}_{k}\in\mathbb{R}^{V}$ and $\mathbf{v}_{k}\in\mathbb{R}^{m}$ of matrices $\mathbf{U}$ and $\mathbf{V}$ the "word analyzer vectors" and "context analyzer vectors," respectively. These vectors represent the alignment of words and contexts with each concept: for a word $z$ and concept $k$, the sign of $\mathbf{u}_{k}[z]$ indicates association ($\mathbf{u}_{k}[z]>0$) or opposition ($\mathbf{u}_{k}[z]<0$) to the concept, while its magnitude quantifies the strength of this relationship. Words for which $|\bm{u}_{k}[z]|$ is small are neutral to this concept. The same interpretation applies to context components $\mathbf{v}_{k}[j]$. We can explicitly represent concepts in the embedding space by defining $d$-dimensional concept representations $\bm{u}^{\rm{d}}_{k}={\bm{W}}^{\top}\bm{u}_{k}$ and ${\bm{v}}^{\rm{d}}_{k}={\bm{H}}{\bm{v}}_{k}$ for $k\in[r]$ as projections onto the word and context spaces, respectively. These can also be viewed as weighted averages of word or context embeddings (similar to classical semantic axes) where the entries of $\bm{u}_{k}$ (or ${\bm{v}}_{k}$) reflect each word’s (or context’s) relevance to the concept.
For an initial confirmation of this interpretation of SVD factors, we revisit the dataset with syntax "The organism that [attribute] is [subject]", [^3] For visual clarity, this design isolates semantic concept extraction by minimizing grammatical influences. shown in Fig. 2C. Fig. 1A shows the columns of $\bm{U}$ and $\bm{V}$}, that is, the word and context analyzer vectors. Inspecting them reveals how they can capture semantic information through their sign patterns: For example, the first column/dimension is such that plant-related words and contexts have positive entries, while animal-related ones have negative entries, indicating that this concept represents the semantic distinction between plants and animals.
However, the figure also reveals that not all individual concepts correspond to human-interpretable categories. Instead, we posit that interpretable semantics emerge from specific combinations of several concepts. We formalize this in the next section and demonstrate in Sec. 4 that it yields interpretable semantics on larger-scale datasets.
3.2 Interpretable Semantics via Orthant-based Clustering
We extend the sign-based interpretation of individual concepts to combinations of multiple concepts. Our key claim is that combining a few top concepts according to their sign patterns, a method we call orthant-based clustering, yields interpretable semantic categories.
For a selected set $\mathcal{K}=\{k_{1},\ldots,k_{p}\}$ of $p\leq r$ concepts, there are $2^{p}$ possible sign configurations $C=[c_{k_{1}},\ldots,c_{k_{p}}]\in\{\pm 1\}^{p}$, where $c_{k_{i}}\in\{\pm 1\}$ indicates whether dimension $k_{i}$ contributes positively or negatively to a semantic category. Each configuration defines a potential semantic category with member words and their degrees of association with that category, formalized as follows:
Definition 1.
For $C=[c_{k_{1}},\ldots,c_{k_{p}}]$ where each $c_{k_{i}}=\pm 1$ for concept dimensions $k_{i}\in[r]$:
$\bullet$ A word $z$ is a member of $C$ if and only if $\texttt{sign}(\bm{u}_{k_{i}}[z])=c_{k_{i}}$ for all $k_{i},i\in[p]$.
$\bullet$ The typicality of a member $z$ of $C$ is defined as: ${\rm{Typicality}}(z;C)=\sum\nolimits_{i\in[p]}\big|\bm{u}_{k_{i}}[z]\big|\,.$
Analogous definitions apply to contexts.
Geometrically, configuration $C$ corresponds to an orthant in the $p$-dimensional subspace spanned by selected concepts. Membership indicates which word embeddings lie within this orthant, while typicality measures the $\ell_{1}$-norm of a word’s projection onto the selected concept dimensions.